Ail is paid $1.25 for every tree he plants.He also receives a bonus of $5 for every 100 trees he plants.How many trees must Ail plant to earn $690?
1
Expert's answer
2011-08-09T07:45:00-0400
Let's estimate how much approximately can be bought for $690.
$690/ 1.25 = 552 trees.
in 552 there are 5 hundreds. For each hundred you receive $5, hence 5*$5 = $25.
Denote the number of trees as n.
Thus
$1.25 * n& - $25 =$690
n = (690 + 25)/1.25 = 572.
In 572 there are still 5 hundreds, so the answer 572 is correct.